OpenStudy (anonymous):

my last question and i believe its d. Solve the equation for y: 6x + 11y = 17 A. y = -6x + 17 B. y = -6x/11 + 17 C. 11y = -23x D. y = -6x/11 + 17/11

OpenStudy (anonymous):

6x + 11y = 17 Subtract 6x to both sides: 11y = -6x + 17 Divide 11 to both sides: y = -6/11x + 17/11 Nice work \(\huge\color{lime} \checkmark\)